OpenStudy (tylermckinney16):

Jake has $100 income to spend on CD's (C) and books (B). A CD costs $10 and a book costs $20. Write the equation of the budget line for this situation in slope intercept form (c should be your x variable and b should be your y variable). Show your work.

OpenStudy (emogirlatemoooocow):

like y=mx+b?

OpenStudy (tylermckinney16):

I cant message back cause it says i can only message you if you fan me lol

OpenStudy (emogirlatemoooocow):

oh yeah i for got about that xD

OpenStudy (tylermckinney16):

lol XD and idk what this question means im dumb af lol XD

OpenStudy (tylermckinney16):

Im joking lol and you should fan me lol.

OpenStudy (emogirlatemoooocow):

Like im Confused As well . If its not y=mx+b then i dont know ..

OpenStudy (tylermckinney16):

I know I been trying to figure it out for a wile lol.

OpenStudy (emogirlatemoooocow):

@Elsa213 Do you Understand this ?

OpenStudy (austen831):

\[100\ge10c+20b\]

OpenStudy (emogirlatemoooocow):

Ahhh so it wasss y=mx+ b ;-;

OpenStudy (tylermckinney16):

^^^^ you had it after all lol XD

OpenStudy (emogirlatemoooocow):

Yeah xD

OpenStudy (emogirlatemoooocow):

@Austen831 thanks :3

OpenStudy (austen831):

no problem